# Drivel-ology: Challenging LLMs with Interpreting Nonsense with Depth

### v1

@@ A.1 Overview of the Annotation Process
여자친구:나 살찐 거 같아?남자친구:넌 살이 문제가 아니야。 "Girlfriend: Do you think I gained weight? Boyfriend: Your problem isn’t your weight." inversion
집에 불이 났다. 온 가족이 당황해서 소리친다. 아버지 : "야, 119가 몇 번이야? 119가 몇 번이냐고 !!! " 아들 : "아버지, 이럴 때일수록 침착하셔야 돼요. 제가 114에 전화해서 물어볼게요    " The house caught fire. The whole family was panicking and shouting. Father: "Hey, what’s the number for 119? What’s the number for 119!!!" Son: "Dad, you need to stay calm in situations like this. I’ll call 114 and ask." misdirection, switchbait
Table 3: Representative examples of Drivelology. The DrivelHub dataset is available on the HuggingFace Hub at https://huggingface.co/datasets/extraordinarylab/drivel-hub.
@@ A.2 Dataset Distribution
Table 4 presents the language distribution of samples in the DrivelHub dataset. As shown, the dataset is skewed toward Mandarin, which accounts for 277 out of the total Drivelology samples. In contrast, other languages such as Japanese and Korean are present only in limited quantities. To characterise the distribution and overlap of annotation categories in our dataset, we present an UpSet plot Lex et al. (2014) is presented in Figure  5, summarising intersections among the five Drivelology categories.
@@ B.1 Experiment Prompts
To ensure reproducibility and transparency, we provide the exact prompts used in each of our experimental tasks. These prompts were carefully designed to probe different aspects of Drivelology comprehension and generation across various LLMs. Below, we detail the prompts for each task: Drivelology Detection (Figure 7), Drivelology Tagging (Figure 8), Narrative Writing (Figure 9), Narrative Selection (Figure 10 for Easy and Figure 11 for Hard).
@@ Appendix C Drivelology Generation
We designed a two experiment settings: (1) Minimal Guidance and (2) With Category Definitions. In the Minimal Guidance setting, the model received only a brief introduction to Drivelology, without any example texts or category definitions. In the Category Definitions setting, the model was provided with detailed definitions of the five Drivelology categories (see §3.1). For each stage, we tested three prompting strategies: zero-shot, one-shot, and five-shot.
@@ C.1 Findings
Our investigation into GPT-4’s generative capabilities reveals a significant gap between mimicking linguistic forms and achieving genuine pragmatic depth. The model’s performance varied with the level of guidance, but it consistently struggled to capture the essential qualities of Drivelology.
@@ D.2 Developing Metrics for Drivelology Generation
Our current study focuses primarily on the understanding and reasoning abilities of LLMs rather than their capacity for generation. A significant area for future work is to establish a comprehensive framework for evaluating generated Drivelology. A key limitation to address is the absence of metrics capable of quantifying the qualities that make a text Drivelological.
Simply measuring fluency or grammatical correctness is insufficient. A robust evaluation would require developing novel metrics to assess specific aspects of a generated Drivelology text, such as: (1) Entertainability: How humorous, witty, or engaging is the output? (2) Relevance: Given a specific topic or context, how relevant is the generated Drivelology? (3) Cohesion and Paradoxical Depth: How well does the output maintain surface-level coherence while simultaneously embedding a meaningful, non-obvious contradiction or twist?
Developing such a framework is a challenging but essential next step. It would not only enable a more rigorous assessment of generative models but also provide clearer objectives for training models to produce high-quality, creative, and contextually aware Drivelology.

### v2

@@ A.1 Overview of the Annotation Process
學生:老師,我媽要我問一下我的成績出來了嗎?老師:你等一下。學生:好的。老師:09 55 34 20 47。學生:打不通。老師:這是成績。 Student: Teacher, my mom asked me to check if my grades are out yet? Teacher: Just a moment. Student: Okay. Teacher: 09 55 34 20 47. Student: Can’t get through. Teacher: That’s your grade. misdirection
我:今年過年我要帶女朋友回去喔。老媽:幾歲,哪裡人?我:到時候你們自己問他,他很溫柔可愛體貼,沒什麼缺點。老媽:你就是他最大的缺點。 Me: This year during Lunar New Year, I’m bringing my girlfriend home. Mom: How old is she? Where is she from? Me: You can ask her yourself then. She’s gentle, cute, thoughtful—she doesn’t have many flaws. Mom: You’re her biggest flaw. misdirection
@@
여자친구:나 살찐 거 같아?남자친구:넌 살이 문제가 아니야。 Girlfriend: Do you think I gained weight? Boyfriend: Your problem isn’t your weight. inversion
집에 불이 났다. 온 가족이 당황해서 소리친다. 아버지 : 야, 119가 몇 번이야? 119가 몇 번이냐고 !!!아들: 아버지, 이럴 때일수록 침착하셔야 돼요. 제가 114에 전화해서 물어볼게요 The house caught fire. The whole family was panicking and shouting. Father: Hey, what’s the number for 119? What’s the number for 119!!! Son: Dad, you need to stay calm in situations like this. I’ll call 114 and ask. misdirection, switchbait
Quel est le coquillage le plus léger? La palourde. What is the lightest shell? The clam. wordplay
Une vague amoureuse du vent lui demande : Est-ce que tu peux me faire une petite bise aujourd’hui? A wave, in love with the wind, asks: “Can you give me a little kiss today?” wordplay
Que horrible cuando tu mamá te da instrucciones y tú estás medio dormida, entonces no te acuerdas si tenías que lavar la basura, colgar al perro o sacar a pasear la ropa. How awful when your mom gives you instructions while you’re half asleep, so you can’t remember whether you were supposed to wash the trash, walk the dog, or take the clothes out for a walk. misdirection
C’est l’histoire de deux pommes de terre. Une d’elles se fait écraser et l’autre s’écrie: Oh purée ! It’s the story of two potatoes. One of them gets crushed, and the other exclaims: Oh mashed potatoes! wordplay
Pourquoi est-ce que Hulk a un beau jardin? Parce qu’il a la main verte. Why does Hulk have a beautiful garden? Because he has a green thumb. switchbait, wordplay
Que fait un employé de chez Sephora à sa pause clope ? Il parfumer. What does a Sephora employee do during their cigarette break? They perfume. switchbait, wordplay
Qu’est ce qu’une lampe moche ? Un LEDron. What do you call an ugly lamp? A LED-boring. wordplay
Pourquoi est ce que Potter est triste ? Parce que personne Harry à sa blague. Why is Potter sad? Because no one Harry gets his joke. switchbait, wordplay
Table 3: Representative examples of Drivelology.
@@ A.2 Dataset Distribution
Table 4 presents the language distribution of samples in the DrivelHub dataset. As shown, the dataset is skewed toward Mandarin, which accounts for 277 out of the total Drivelology samples. In contrast, other languages such as Japanese and Korean are present only in limited quantities. To characterise the distribution and overlap of annotation categories in our dataset, we present an UpSet plot Lex et al. (2014) in Figure  5, summarising intersections among the five Drivelology categories.
@@ B.1 Experiment Prompts
To ensure reproducibility and transparency, we provide the exact prompts used in each of our experimental tasks. These prompts were carefully designed to probe different aspects of Drivelology comprehension and generation across various LLMs. Below, we detail the prompts for each task: Drivelology Detection (Figure 7), Drivelology Tagging (Figure 8), Narrative Writing (Figure 9 for generation and Figure 10 for evaluation), Narrative Selection (Figure 11 for Easy and Figure 12 for Hard).
@@ Appendix C Drivelology Generation
We designed two experimental settings: (1) Minimal Guidance and (2) Guidance With Category Definitions. In the Minimal Guidance setting, the model received only a brief introduction to Drivelology, without any example texts or category definitions. In the Category Definitions setting, the model was provided with detailed definitions of the five Drivelology categories (see §3.1). For each stage, we tested three prompting strategies: zero-shot, one-shot, and five-shot.
@@ C.1 Findings
Our investigation into GPT-4’s generative capabilities reveals a significant gap between mimicking linguistic forms and achieving genuine pragmatic depth.
@@ D.2 Developing Metrics for Drivelology Generation
Our current study focuses primarily on the understanding and reasoning abilities of LLMs rather than their capacity for generation. A significant area for future work is to establish a comprehensive framework for evaluating generated Drivelology. A key limitation to address is the absence of metrics capable of quantifying the qualities that make a text Drivelological. Simply measuring fluency or grammatical correctness is insufficient. A robust evaluation would require developing novel metrics to assess specific aspects of a generated Drivelology text, such as: (1) Entertainability: How humorous, witty, or engaging is the output? (2) Cohesion and Paradoxical Depth: How well does the output maintain surface-level coherence while simultaneously embedding a meaningful, non-obvious contradiction or twist? (3) Originality: How surprising or non-formulaic is the output? Does it avoid simply rehashing common Drivelological text or existing templates? (4) Cultural Resonance: How well does the output tap into shared cultural knowledge, social scripts, or contemporary memes to create its meaning?
Furthermore, a robust framework must test for controllable generation – the ability to create Drivelology that meets specific constraints, like producing an “inversion” about “technology.” Success here would be a strong signal of true comprehension. While developing this framework is challenging, it is essential for two reasons: it would allow for more rigorous model assessment and provide clearer targets for training. This creates a powerful feedback loop where better evaluation drives better generation, which in turn deepens the model’s core reasoning, ultimately leading to LLMs that can truly master “nonsense with depth.”
@@
Task:
Evaluate how accurately the candidate narrative matches the given reference narrative.
Use a scale from 1 to 5, where 1 indicates the least accuracy and 5 indicates the highest accuracy.
- Candidate Narrative: {candidate}
- Reference Narrative: {reference}
Output Format:
Please provide the output in JSON format with the following key:
- score: The score indicating the level of matching, ranging from 1 to 5.
Figure 10: Prompts for evaluating Narrative Writing task.
Prompt1: